OpenStudy (anonymous):

Could someone check my answer?? Find the derivative of the function by using the definition.

OpenStudy (anonymous):

\[y = \frac{ 3 }{x+2} ; x = 4\]

OpenStudy (anonymous):

I came up with 0

OpenStudy (anonymous):

y' = -3(x+2)^-2

OpenStudy (anonymous):

how did you get that if you dont mind me asking?